S= 2B+PH solve for H
Found 2 solutions by rfer, richwmiller:
Answer by rfer(16322)

s=2b+ph
s-2b=ph
(s-2b)/p=h

Answer by richwmiller(17219)

we want to leave just h on the right
subtract 2b from both sides
s-2b=ph
divide both sides by p
(s-2b)/p=h